If the diagonal of a rectangle is 25cms
and the short sides of the rectangle are 9cms
How do you find the length of the long side???

Answer by checkley77(12844)

The diagonal of a rectancle forms a right triangle whose sides are a^2+b^2=c^2.
9^2+x^2=25^2
81+x^2=625
x^2=625-81
x^2=544
x=sqrt544
x=23.3238 ans. for the long side.